HETAV PATEL

Grade 11,

You have discovered a new radioisotope 19X with half life of 14.2 days. How long would it take a solution containing 42000 dpm to decay to 500 dpm?

You have discovered a new radioisotope 19X with half life of 14.2 days. How long would it take a solution containing 42000 dpm to decay to 500 dpm?

Grade:12th pass

1 Answers

noogler
8 years ago
as it is a radioactive substance it will undergo first order 
use t1/2=0.693/k...........................find k
now t=(2.303/k)log(a/a-x)...........................find t
t is required time 
solve it........................................................................
APPROVE IT HELPED U TO SOLVE THIS PROBLEM